OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, -6) and is perpendicular to y = 3x - 6. y =1/3 x - 8 y = -3x + 12 y = 3x - 24 y = -1/3x - 4

OpenStudy (anonymous):

y=-1/3x-4

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

If two lines are perpendicular, then their slopes are related m1=-1/m2. Our slope is perpendicular to slope 3, so the line has a slope -1/3. If the two lines are to intersect, then the point (6,-6) must satisfy the linear equation. -6 = (-1/3)(6)+b ===> b=-4 y=-1/3 x -4

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome =)